OpenStudy (anonymous):

two number cubes were rolled 150 times. Find the experimental probability that a sum of 6 or greater appeared 85 times anybody?

OpenStudy (shubhamsrg):

this is a set of bernoulli's trials, heard about them ?

OpenStudy (anonymous):

set up an paporsas so it like 6/85 x/150 so 150x6= then divide by 85

OpenStudy (anonymous):

thankyou. :)

OpenStudy (anonymous):

\[(150 C 85)\left( \frac{ 26 }{ 36 } \right)^{85}\left( \frac{ 10 }{ 36 } \right)^{65}\]

OpenStudy (anonymous):

The 26/36 comes from the individual roll probability of getting a 6 or more, consequently, the 10/36 is the individual roll probability of rolling a 5 or less. The "150 combination 85" accounts for all variations or orders of individual outcomes within the 150 rolls.

OpenStudy (anonymous):

All good now, @erikalove951 ?

OpenStudy (anonymous):

yes, thankyou @tcarroll010

OpenStudy (anonymous):

:)